OpenStudy (anonymous):

What are the equations for the asymptotes of this hyperbola? y^2/36-x^2/121=1 A. y=√(157/5)x, y= -√(157/5)x B. y=(6/11)x, y=-(6/11)x C. y=(11/6)x, y=-(11/6)x D. y=(√85/5)x, y=-(√85/5)x

OpenStudy (anonymous):

y = -(6 x)/11 | y = (6 x)/11

OpenStudy (anonymous):

B

OpenStudy (amistre64):

should be y/x sqrt(36/121) = 6/11